# Incentive-Compatible Vertiport Reservation in Advanced Air Mobility: An Auction-Based Approach

### v2

@@ 2.2 Problem Formulation
Given an allocation ${x}$, let $S(r,t,{x})\in\mathbb{Z}_{+}$ denote the state of vertiport – defined to be the number of aircraft occupying the parking spots at vertiport $r\in R$ at time $t\in[H]$. For every $r\in R$, the initial occupation $S(r,1,{x})$ is defined as
$S(r,1,x):=\sum_{i\in F}\sum_{j\in A_{i}}\mathbf{1}(r^{\textsf{orig}}_{i,j}=r).$
For concise notation, we shall denote $S(r,1,x)$ by $\bar{S}(r)$ for every $r\in R$ as it does not depend on $x$. Naturally, it must hold that, for every $r\in R,t\in\{2,\ldots,H\}$,
@@
    Each aircraft is allocated at most one route. That is, for every $i\in F,j\in A_{i}$,
    $\sum_{k\in m_{i,j}}{x}_{i,j,k}\leq 1.$
@@
Consequently, the set of feasible allocation is defined as
@@
The social welfare of any allocation $x\in X$ is defined as follows.
@@
where $(i)$ $\rho_{i}\in\mathbb{R}_{+}$ is the weight factor specifying the relative importance of different FOs [^5] The coefficient $\rho_{i}$ is reminiscent of remote city opportunity factor, are used in (Dixit et al., 2023)., $(ii)$ $C_{r,t}:\mathbb{Z}_{+}\rightarrow\mathbb{R}_{+}$ with $C_{r,t}(0)=0$ is discrete convex [^6] Based on (Murota, 2015), a function $f:\mathbb{Z}\rightarrow\mathbb{R}$ is discrete convex if $f(x+1)-f(x)\geq f(x)-f(x-1),\ \forall x\in\mathbb{Z}$. to capture increasing marginal cost of congestion, and $(iii)~{}\lambda\in\mathbb{R}_{+}$ is the ratio between the congestion cost and the cumulative weighted valuations of FOs. Furthermore, we define an optimal allocation as
@@
The social welfare objective (3) captures three main desiderata: efficiency, fairness, and safety. The objective (3) incorporates efficiency through additive valuations of FOs. Additionally, it incorporates the proportional fairness criterion by assigning different weights to the valuations of different FOs, denoted by $(\rho_{i})_{i\in F}$. Well-constructed weights can prevent larger FOs from monopolizing the resources. Finally, it encompasses safety considerations in two ways: first, through capacity constraints; and second, by introducing a congestion-dependent term in (3) that penalizes vertiports when the number of aircraft increases. With these three considerations, the definition of social welfare aligns closely with that presented in (Dixit et al., 2023).
We assume the SP does not have access to the true valuations $V$, as it is private information. Instead, the SP must use bids $B$ reported by the FOs to allocate the aircraft to vertiports through an auction mechanism – vertiport reservation mechanism. More formally, given a bid profile $B$, the SP uses a mechanism $\bar{M}=(\bar{{x}},(\bar{p}_{i})_{i\in F})$, where for a given bid profile $B$, $(i)$ $\bar{{x}}(B)\in X$ is the allocation proposed by the mechanism; and $(ii)$ $\bar{p}_{i}(B)\in\mathbb{R}$ denotes the payment charged to FO $i\in F$. Under the mechanism $\bar{M}$, the utility derived by any FO $i\in F$ is
$U_{i}(B;\bar{M}):=\sum_{j\in A_{i}}\sum_{k\in m_{i,j}}v_{i,j,k}\textbf{1}(\bar{{x}}_{i,j,k}(B))-\bar{p}_{i}(B).$ (5)
@@
    Incentive Compatibility (IC): Bidding truthfully is each FO’s (weakly) dominant strategy, i.e., for every $i\in F$, $B_{-i}\in\mathbb{R}_{+}^{L_{-i}}$,
    $V_{i}\in\underset{B_{i}\in\mathbb{R}_{+}^{L_{i}}}{\arg\max}~{}U_{i}(B_{i},B_{-i};\bar{M}).$
@@
    $U_{i}(V_{i},B_{-i};\bar{M})\geq 0,\quad\forall\ B_{-i}\in\mathbb{R}_{+}^{L_{-i}}.$
@@ 3 Mechanism Design
We present an auction mechanism that satisfies (D1)-(D3) in Section 3.1, and prove its theoretical properties in Section 3.2. Moreover, we present an approach for computing the auction mechanism in Section 4.
@@ 3.1 Mechanism
Allocation Function: Given a bid profile $B\in\mathbb{R}^{L}_{+}$, the allocation is obtained by
@@
Payment Function: We first define a function $\theta:F\times\mathbb{R}_{+}^{L}\rightarrow\mathbb{R}_{+}^{L}$ such that for any $i\in F,j\in A_{i},k\in m_{i,j},$
$\forall~{}\ell\in F,B\in\mathbb{R}_{+}^{L},\quad\theta_{i,j,k}(\ell,B)=\begin{cases}b_{i,j,k},&\text{if}\ i\neq\ell,\\ 0,&\text{if}\ i=\ell.\end{cases}$ (7)
@@
where for every $i\in F$, ${x}\in X$, and $B\in\mathbb{R}_{+}^{L}$,
@@
There are some notable differences between the VCG payment and (8). First, since our problem is an “exchange problem” and not the typical “assignment problem”, we need to be cognizant of the physical resources occupied by the aircraft. However, this would require us to enumerate all the feasible combinations if we were to directly implement VCG mechanisms. To overcome the problem of enumerating all feasible solutions while computing payments, we adopt a novel approach of “pseudo-bids”, where while computing the payments, each non-participating aircraft is considered to be using a bid of 00, as formally described in (7).
Second, since the objective function (3) is not the summation of the participants’ valuations, the typical VCG auction is not directly applicable. Instead, we follow (Dixit et al., 2023; Nisan et al., 2007) to devise the payment rule for any $i\in F$ and $b\in\mathbb{R}_{+}^{L_{i}}$.
@@ 3.2 Theoretical Analysis
The proposed mechanism $\bar{M}=(\bar{x},\bar{p})$ is IC, IR, and SWM (i.e. satisfied (D1)-(D3)).
@@ 4 Optimization Algorithm
In this section, we formulate (6) as a mixed binary linear program (MBLP), as shown in (16). We derive this in three steps. First, in Section 4.1, we construct a time-extended flow network, where vertices are vertiport-time and aircraft-time pairs with edges capturing capacity constraints and route allocation. Then, using binary variables ($\delta_{i,j,\tau}$ as formally defined later in (12d) and (12e)) to ensure that each aircraft is allocated one route, we formulate a mixed integer linear program (MILP (12)) in Section 4.2. This MILP has fewer binary variables than (6) when the number of unique departure times for any aircraft is less than the size of its menu. Finally, in Section 4.3, we show that the total unimodularity of the constraint matrix ($\bar{I}_{\star}$ in (12b)) guarantees that all flows are integral for each binary variable assignment, so we can drop the integrality constraint (12f) and get the final MBLP formulation (16).
@@ 4.1 Auxiliary Graph
      We consider three replica for each vertiport $r\in R$ at time $t\in[H]$, denoted as $\bar{\nu}(r,t),\bar{\nu}^{\textsf{arr}}(r,t),$ and $\bar{\nu}^{\textsf{dep}}(r,t)$.
@@
      For each $i\in F,j\in A_{i},$ we consider one vertex corresponding to all routes that have the same departure time. More formally, for every $i\in F,j\in A_{i},$ define $T^{\textsf{dep}}_{i,j}:=\cup_{k\in m_{i,j}}\left\{t^{\textsf{dep}}_{i,j,k}\right\},$ to be the set of unique departure times amongst all routes. We consider one vertex corresponding to each $i\in F,j\in A_{i},$ and $\tau\in T^{\textsf{dep}}_{i,j}$, denoted as $\bar{\nu}({i,j,\tau})$.
@@
      $\bar{\nu}^{\textsf{source}}$ and $\bar{\nu}^{\textsf{sink}}$ denote the source and sink in the flow network (to be described shortly).
@@
      For every $r\in R,t\in[H-1]$, we consider $\textsf{park}(r,t)$ number of edges connecting $\bar{\nu}(r,t)$ and $\bar{\nu}(r,t+1)$. We denote this set by $\bar{E}_{3,r,t}$. For any $q\in[\textsf{park}(r,t)]$, we denote the $q-$th edge in $\bar{E}_{3,r,t}$ by $e_{q,r,t}$ which is defined to be the following tuple
      $e_{q,r,t}=(\bar{\nu}(r,t),\bar{\nu}(r,t+1),\overline{\textbf{c}}=1,\underline{\textbf{c}}=0,\bar{\textbf{w}}=-\lambda(C_{r,t}(q)-C_{r,t}(q-1))).$
@@
      For every $r\in R$, we consider $\textsf{park}(r,H)$ number of edges connecting $\bar{\nu}(r,H)$ and $\bar{\nu}^{\textsf{sink}}$. We denote these edges by $\bar{E}_{8,r}$. For any $q\in[\textsf{park}(r,t)]$, we denote the $q-$th edge in $\bar{E}_{8,r}$ by $e_{q,r,H}$ which is defined to be the following tuple
      $e_{q,r,H}=(\bar{\nu}(r,H),\bar{\nu}^{\textsf{sink}},\overline{\textbf{c}}=1,\underline{\textbf{c}}=0,\bar{\textbf{w}}=-\lambda(C_{r,H}(q)-C_{r,H}(q-1))).$
@@ 4.2 Mixed Integer Linear Program Formulation
For the construction of auxiliary graph presented in Section 4.1, we concatenate the weight, upper capacity bound, and lower capacity bound of each edge as $\overline{\mathbf{W}}\in\mathbb{R}^{|\bar{E}|}$, $\mathbf{\overline{C}}(\delta)\in\mathbb{Z}_{+}^{|\bar{E}|}$, and $\mathbf{\underline{C}}(\delta)\in\mathbb{Z}_{+}^{|\bar{E}|}$, respectively. Define an incidence matrix of the graph $\bar{G}$ as $\bar{I}\in\{-1,0,1\}^{|\bar{V}|\times|\bar{E}|}$, where
@@
Defining a truncated incidence matrix $\bar{I}_{\star}$ obtained from $\bar{I}$ by removing rows corresponding to $\bar{\nu}^{\textsf{source}}$ and $\bar{\nu}^{\textsf{sink}}$, we consider the following optimization formulation (which will be shown to be crucial in computation of (6)).
@@
$\ \ \ \ \ \ \ {\mathbf{A}_{e_{q+1,r,t}}\!\leq\!\mathbf{A}_{e_{q,r,t}},\forall r\in R,t\in[H],q\in[\textsf{park}(r,t)-1]}.$ (12g)
@@
The above construction specifies the values of $\textbf{A}_{e}(x)$ for $e\in\bar{E}_{3}\cup\bar{E}_{5}\cup\bar{E}_{8}$. Additionally, by Lemma 4.3, there exists a unique feasible solution $(\textbf{A}(x),\delta(x))$ that satisfied (12b)-(12g) , and we get
@@
$=\sum_{r\in R}\sum_{t=1}^{H-1}\sum_{q=1}^{\textsf{park}(r,t)}\bar{\textbf{w}}_{e_{q,r,t}}\mathbf{A}_{e_{q,r,t}}(x)$
@@
Similarly, we get $\sum_{e\in\bar{E}_{8}}\bar{\textbf{w}}_{e}\mathbf{A}_{e}(x)=-\lambda\sum_{r\in R}C_{r}(S(r,H,x))$. To summarize, we obtain
@@
By $\sum_{e\in\bar{E}_{5}}\textbf{A}_{e}\mathbf{1}((\bar{\nu}({i,j,t^{\textsf{dep}}_{i,j,k}}),\bar{\nu}^{\textsf{arr}}({r^{\textsf{dest}}_{i,j,k},t^{\textsf{arr}}_{i,j,k}}))\in e)=x_{i,j,k}(\textbf{A},\delta)$, we get
$\sum_{i\in F}\sum_{j\in A_{i}}\sum_{k\in m_{i,j}}{x}_{i,j,k}\mathbf{1}(r^{\textsf{dest}}_{i,j,k}=r,t^{\textsf{arr}}_{i,j,k}=t)\leq\textsf{arr}(r,t)$
Note that if $t=1$, the arrival capacity constraints are trivially satisfied since there is no incoming aircraft. Analogously, the flow balance equations at the nodes of the form $\bar{\nu}^{\textsf{dep}}(r,t)$, for some $r\in R,t\in[H]$, ensure that $\sum_{i\in F}\sum_{j\in A_{i}}\sum_{k\in m_{i,j}}{x}_{i,j,k}\mathbf{1}(r^{\textsf{orig}}_{i,j}=r,t^{\textsf{dep}}_{i,j,k}=t)\leq\textsf{dep}(r,t).$ Finally, we can establish $S(r,t,x(\textbf{A},\delta))=\sum_{q=1}^{\textsf{park}(r,t)}\textbf{A}_{e_{q,r,t}}$ through the flow balance equation at $\bar{\nu}(r,t)$ and (1). Since $\sum_{q=1}^{\textsf{park}(r,t)}\textbf{A}_{e_{q,r,t}}\leq\textsf{park}(r,t)$, due to the capacity constraints on the edge $\bar{E}_{3,r,t}$, it holds that $S(r,t,x(\textbf{A},\delta))\leq\textsf{park}(r,t)$. Thus, we conclude that $x(\textbf{A},\delta)\in X$.
@@
Next, we present a technical lemma used in the proof of Proposition 4.2.
Lemma 4.3.
Proof.
First, note that any feasible solution to (12b)-(12g) has the same value of $\textbf{A}_{e}(x)$ for $e\in\bar{E}_{6}\cup\bar{E}_{7}\cup\bar{E}_{9}$ since the lower and upper bound on capacity are the same on these edges by construction. Thus, it is sufficient to show that the values of $\textbf{A}_{e}$ for $e\in\bar{E}_{3}\cup\bar{E}_{5}\cup\bar{E}_{6}\cup\bar{E}_{7}\cup\bar{E}_{8}\cup\bar{E}_{9}$, uniquely determine a feasible solution $(\textbf{A},\delta)$ that satisfies (12b)-(12g). Particularly, we will show that we can uniquely recover the values of $\textbf{A}_{e}$ for $e\in\bar{E}_{1}\cup\bar{E}_{2}\cup\bar{E}_{4}.$
Vertex Incoming Edges Outgoing Edges
$\bar{\nu}^{\textsf{arr}}(r,t)$ ${\bar{E}_{5}}$ $\bar{E}_{1}$
$\bar{\nu}(i,j,\tau)$ $\bar{E}_{4}$ ${\bar{E}_{5}}$
$\bar{\nu}^{\textsf{dep}}(r,t)$ $\bar{E}_{2}$ ${\bar{E}_{4}}$
$\bar{\nu}(i,j,0)$ $\bar{E}_{7}$ $\bar{E}_{9}$
Note that flow balance at nodes of the form $\bar{\nu}^{\textsf{arr}}(r,t)$ will determine the values $\textbf{A}_{e}$ on edge $\bar{E}_{1}$, as we know these values for all edges in the set $\bar{E}_{5}$. Next, flow balance at nodes of the form $\bar{\nu}(i,j,\tau)$ will determine the values $\textbf{A}_{e}$ on edge $\bar{E}_{4}$, as we know these values for edges $\bar{E}_{5}$. This and the capacity constraints on $\bar{E}_{4}$, ensure that we know the value of $\delta$. Next, flow balance at nodes of the form $\bar{\nu}^{\textsf{dep}}(r,t)$ will determine the values $\textbf{A}_{e}$ on edge $\bar{E}_{2}$, as we can uniquely determine these values on $\bar{E}_{4}$. Finally, flow balance at nodes of the form $\bar{\nu}(r,t)$ will determine the values $\textbf{A}_{e}$ on edge $\bar{E}_{2}$, as we can uniquely determine these values on $\bar{E}_{1}\cup\bar{E}_{3}\cup\bar{E}_{6}\cup\bar{E}_{8}\cup\bar{E}_{9}$.

### v3

@@ 2.2 Problem Formulation
Given an allocation ${x}$, let $S(r,t,{x})\in\mathbb{Z}_{+}$ denote the number of aircraft occupying the parking spots at vertiport $r\in R$ at time $t\in[H]$. For every $r\in R$, the initial occupation $S(r,1,{x})$ is
$S(r,1,x)=\sum_{i\in F}\sum_{j\in A_{i}}\mathbf{1}(r^{\textsf{orig}}_{i,j}=r).$
For concise notation, we shall denote $S(r,1,x)$ by $\bar{S}(r)$ for every $r\in R$ since it does not depend on $x$. Naturally, it must hold that, for every $r\in R,t\in\{2,\ldots,H\}$,
@@
    Each aircraft is allocated at most one route. That is, for every $i\in F,j\in A_{i}$, $\sum_{k\in m_{i,j}}{x}_{i,j,k}\leq 1$.
@@
Consequently, we define
@@
Given ${x}\in X$, social welfare is defined as follows.
@@
where $(i)$ $\rho_{i}\in\mathbb{R}_{+}$ is the weight factor specifying the relative importance of different FOs [^5] Similar weight factors, termed as remote city opportunity factor, are used in (Dixit et al., 2023)., $(ii)$ $C_{r,t}:\mathbb{Z}_{+}\rightarrow\mathbb{R}_{+}$ with $C_{r,t}(0)=0$ is discrete convex [^6] Based on (Murota, 2015), a function $f:\mathbb{Z}\rightarrow\mathbb{R}$ is discrete convex if $f(x+1)-f(x)\geq f(x)-f(x-1),\ \forall x\in\mathbb{Z}$. to capture increasing marginal cost of congestion [^7] While we only consider the congestion resulting from parked aircraft, it is straightforward to extend our formulation to arriving and departing aircraft; see Subsection 5.2., and $(iii)~{}\lambda\in\mathbb{R}_{+}$ is the ratio between the congestion cost and the cumulative weighted valuations of FOs. Furthermore, we define an optimal allocation as
@@
The social welfare objective (3) captures three main desiderata: efficiency, fairness, and safety. The objective (3) incorporates efficiency through additive valuations of FOs. Additionally, it incorporates the proportional fairness criterion [^8] We emphasize that the fairness is at the FO-level. by assigning different weights to the valuations of different FOs, denoted by $(\rho_{i})_{i\in F}$. Well-constructed weights can prevent larger FOs from monopolizing the resources; for example, using the logarithm of the number of aircraft as an FO’s weight. Finally, it encompasses safety considerations in two ways: first, through capacity constraints; and second, by introducing a congestion-dependent term in (3) that penalizes vertiports when the number of aircraft increases. With these three considerations, the definition of social welfare aligns closely with that presented in (Dixit et al., 2023).
We assume the SP does not have access to the true valuations $V$, as it is private information. Instead, the SP must use bids $B$ reported by the FOs to allocate the aircraft to vertiports through an auction mechanism. More formally, given a bid profile $B$, the SP uses a mechanism $\bar{M}=(\bar{{x}},(\bar{p}_{i})_{i\in F})$, where for a given bid profile $B$, $(i)$ $\bar{{x}}(B)\in X$ is the allocation proposed by the mechanism; and $(ii)$ $\bar{p}_{i}(B)\in\mathbb{R}$ denotes the payment charged to FO $i\in F$. Under the mechanism $\bar{M}$, the utility derived by any FO $i\in F$ is
$U_{i}(B;\bar{M})=\sum_{j\in A_{i}}\sum_{k\in m_{i,j}}v_{i,j,k}\textbf{1}(\bar{{x}}_{i,j,k}(B))-\bar{p}_{i}(B).$ (5)
@@
    Incentive Compatibility (IC): Bidding truthfully is each FO’s (weakly) dominant strategy, i.e., for every $i\in F$, $B_{-i}\in\mathbb{R}_{+}^{\sum_{\ell\in F\backslash\{i\}}\sum_{j\in A_{\ell}}|m_{\ell,j}|}$,
    $V_{i}\in\underset{B_{i}\in\mathbb{R}_{+}^{\sum_{j\in A_{i}}|m_{i,j}|}}{\arg\max}~{}U_{i}(B_{i},B_{-i};\bar{M}).$
@@
    $U_{i}(V_{i},B_{-i};\bar{M})\geq 0,\quad\forall\ B_{-i}\in\mathbb{R}_{+}^{\sum_{\ell\in F\backslash\{i\}}\sum_{j\in A_{\ell}}|m_{\ell,j}|}.$
@@ 3 Mechanism Design
In this section, we present an auction mechanism that satisfies (D1)-(D3) in Subsection 3.1 and prove its theoretical properties in Subsection 3.2. We defer the optimization algorithm to Section 4.
@@ 3.1 Mechanism
Allocation Function: Given a bid profile $B\in\mathbb{R}^{\sum_{i\in F}\sum_{j\in A_{i}}|m_{i,j}|}_{+}$, the allocation is obtained by
@@
Payment Function: We first define a function $\theta:F\times\mathbb{R}_{+}^{{\sum_{i\in F}\sum_{j\in A_{i}}|m_{i,j}|}}\rightarrow\mathbb{R}_{+}^{{\sum_{i\in F}\sum_{j\in A_{i}}|m_{i,j}|}}$ such that for any $\ell\in F$ and bid $B\in\mathbb{R}_{+}^{{\sum_{i\in F}\sum_{j\in A_{i}}|m_{i,j}|}}$,
$\theta_{i,j,k}(\ell,B)=\begin{cases}b_{i,j,k},&\text{if}\ i\neq\ell,\\ 0,&\text{if}\ i=\ell,\end{cases}\ \forall\ i\in F,j\in A_{i},k\in m_{i,j}.$ (7)
@@
where for every $i\in F$, ${x}\in X$, and $B\in\mathbb{R}_{+}^{{\sum_{i\in F}\sum_{j\in A_{i}}|m_{i,j}|}}$,
@@
There are some notable differences between the VCG payment and (8). First, since our problem is an “exchange problem” and not the typical “assignment problem”, we need to be cognizant of the physical resources occupied by the aircraft of that operator. However, this would require us to enumerate all the feasible combinations if we were to directly implement VCG mechanisms. To overcome the problem of enumerating all feasible solutions while computing payments, we adopt a novel approach of “pseudo-bids”, where while computing the payments, each non-participating aircraft is considered to be using a bid of 00, as formally described in (7).
Second, since the objective function (3) is not the summation of the participants’ valuations, the typical VCG auction is not directly applicable. Instead, we follow (Dixit et al., 2023; Nisan et al., 2007) to devise the payment rule for any $i\in F$ and $b\in\mathbb{R}_{+}^{\sum_{j\in A_{i}}|m_{i,j}|}$.
@@ 3.2 Theoretical Analysis
The proposed mechanism $\bar{M}:=(\bar{x},\bar{p})$, defined by (6) and (8) is IC, IR, and SWM.
@@ 4 Optimization Algorithm
In this section, we formulate (6) as a mixed binary linear program (MBLP), as shown in (16). We derive this in three steps. First, in Subsection 4.1, we construct a time-extended flow network, where vertices are vertiport-time and aircraft-time pairs with edges capturing capacity constraints and route allocation. Then, using binary variables ($\delta_{i,j,\tau}$ as formally defined later in (12d) and (12e)) to ensure that each aircraft is allocated one route, we formulate a mixed integer linear program (MILP (12)) in Subsection 4.2. This MILP has fewer binary variables than (6) when the number of unique departure times for any aircraft is less than the size of its menu. Finally, in Subsection 4.3, we show that the total unimodularity of the constraint matrix ($\bar{I}_{\star}$ in (12b)) guarantees that all flows are integral for each binary variable assignment, so we can drop the integrality constraint (12f) and get the final MBLP formulation (16).
@@ 4.1 Auxiliary Graph
      We consider three replica for each vertiport $r\in R$ at time $t\in[H]$, denoted as $\bar{\nu}(r,t),\bar{\nu}^{\textsf{arr}}(r,t),$ and $\bar{\nu}^{\textsf{dep}}(r,t)$. These vertices, along with $\bar{E}_{1}$, $\bar{E}_{2}$, $\bar{E}_{3}$, and $\bar{E}_{8}$ defined later, embed capacity constraints and congestion costs into the graph structure.
@@
      For each $i\in F,j\in A_{i},$ we consider one vertex corresponding to all routes that have the same departure time. More formally, for every $i\in F,j\in A_{i},$ define $T^{\textsf{dep}}_{i,j}:=\cup_{k\in m_{i,j}}\left\{t^{\textsf{dep}}_{i,j,k}\right\},$ to be the set of unique departure times amongst all routes. We consider one vertex corresponding to each $i\in F,j\in A_{i},$ and $\tau\in T^{\textsf{dep}}_{i,j}$, denoted as $\bar{\nu}({i,j,\tau})$, which, along with $\bar{E}_{4}$, $\bar{E}_{5}$, $\bar{E}_{7}$, and $\bar{E}_{9}$ defined later, embeds the route choice of the aircraft.
@@
      $\bar{\nu}^{\textsf{source}}$ and $\bar{\nu}^{\textsf{sink}}$ denote the source and sink in the flow network (to be described shortly). These vertices, along with $\bar{E}_{6}$, $\bar{E}_{7}$, and $\bar{E}_{8}$, ensure flow conservation of the parking aircraft.
@@
      For every $r\in R,t\in[H-1]$, we consider $\textsf{park}(r,t)$ edges connecting $\bar{\nu}(r,t)$ and $\bar{\nu}(r,t+1)$. We denote this set by $\bar{E}_{3,r,t}$. For any $q\in[\textsf{park}(r,t)]$, we denote the weight of the $q-$th edge in $\bar{E}_{3,r,t}$ by $\bar{\textbf{w}}_{q,r,t}$, and upper and lower capacity by $\overline{\textbf{c}}_{q,r,t}$ and $\underline{\textbf{c}}_{q,r,t}$, respectively. For any $r\in R,q\in[\textsf{park}(r,t)]$, $\bar{\textbf{w}}_{q,r,t}=-\lambda(C_{r,t}(q)-C_{r,t}(q-1))$, $\overline{\textbf{c}}_{q,r,t}=1$, and $\underline{\textbf{c}}_{q,r,t}=0$.
@@
      For every $r\in R$, we consider $\textsf{park}(r,H)$ edges connecting $\bar{\nu}(r,H)$ and $\bar{\nu}^{\textsf{sink}}$. We denote these edges by $\bar{E}_{8,r}$. For any $q\in[\textsf{park}(r,H)]$, we denote the weight of the $q-$th edge in $\bar{E}_{8,r}$ by $\bar{\textbf{w}}_{q,r,H}$, and upper and lower capacity by $\overline{\textbf{c}}_{q,r,H}$ and $\underline{\textbf{c}}_{q,r,H}$ respectively. For any $r\in R,q\in[\textsf{park}(r,H)]$, $\bar{\textbf{w}}_{q,r,H}=-\lambda(C_{r,H}(q)-C_{r,H}(q-1)),\overline{\textbf{c}}_{q,r,H}=1$, and $\underline{\textbf{c}}_{q,r,H}=0$.
@@ 4.2 Mixed Integer Linear Program Formulation
We concatenate the weight, upper capacity bound, and lower capacity bound of each edge as $\overline{\mathbf{W}}\in\mathbb{R}^{|\bar{E}|}$, $\mathbf{\overline{C}}\in\mathbb{Z}_{+}^{|\bar{E}|}$, and $\mathbf{\underline{C}}\in\mathbb{Z}_{+}^{|\bar{E}|}$, respectively. Define an incidence matrix of the graph $\bar{G}$ as $\bar{I}\in\{-1,0,1\}^{|\bar{V}|\times|\bar{E}|}$, where
@@
Defining a truncated incidence matrix $\bar{I}_{\star}$ obtained from $\bar{I}$ by removing rows corresponding to $\bar{\nu}^{\textsf{source}}$ and $\bar{\nu}^{\textsf{sink}}$, we have the following optimization problem.
@@
$\ \ \ \ \ \ \ \mathbf{A}_{q+1,r,t}\!\leq\!\mathbf{A}_{q,r,t},\forall r\in R,t\in[H],q\in[\textsf{park}(r,t)-1].$ (12g)
@@
The above construction specifies the values of $\textbf{A}_{e}(x)$ for $e\in\bar{E}_{3}\cup\bar{E}_{5}\cup\bar{E}_{8}$. Additionally, by Lemma 4.2, there exists a unique feasible solution $(\textbf{A}(x),\delta(x))$, and we get
@@
$=\sum_{r\in R}\sum_{t=1}^{H-1}\sum_{q=1}^{\textsf{park}(r,t)}\bar{\textbf{w}}_{q,r,t}\mathbf{A}_{q,r,t}(x)$
@@
Similarly, we get $\sum_{e\in\bar{E}_{8}}\bar{\textbf{w}}_{e}\mathbf{A}_{e}(x)=-\lambda\sum_{r\in R}C_{r}(S(r,H,x))$.
To summarize, we obtain
@@
By $\sum_{e\in\bar{E}_{5}}\textbf{A}_{e}\mathbf{1}((\bar{\nu}({i,j,t^{\textsf{dep}}_{i,j,k}}),\bar{\nu}^{\textsf{arr}}({r^{\textsf{dest}}_{i,j,k},t^{\textsf{arr}}_{i,j,k}}))\!\!\in\!\!e)\!\!=\!\!x_{i,j,k}(\textbf{A},\delta)$, we get $\sum_{i\in F}\sum_{j\in A_{i}}\sum_{k\in m_{i,j}}{x}_{i,j,k}\mathbf{1}(r^{\textsf{dest}}_{i,j,k}=r,t^{\textsf{arr}}_{i,j,k}=t)\!\!\leq\!\!\textsf{arr}(r,t)$ [^10] When $t=1$, the arrival capacity constraints are trivially satisfied since there is no incoming aircraft.. Analogously, the flow balance equations at the nodes of the form $\bar{\nu}^{\textsf{dep}}(r,t)$, for some $r\in R,t\in[H]$, ensure that $\sum_{i\in F}\sum_{j\in A_{i}}\sum_{k\in m_{i,j}}{x}_{i,j,k}\mathbf{1}(r^{\textsf{orig}}_{i,j}=r,t^{\textsf{dep}}_{i,j,k}=t)\leq\textsf{dep}(r,t).$ Finally, we can establish $S(r,t,x(\textbf{A},\delta))=\sum_{q=1}^{\textsf{park}(r,t)}\textbf{A}_{q,r,t}$ through the flow balance equation at $\bar{\nu}(r,t)$ and (1). Since $\sum_{q=1}^{\textsf{park}(r,t)}\textbf{A}_{q,r,t}\leq\textsf{park}(r,t)$, due to the capacity constraints on the edge $\bar{E}_{3,r,t}$, it holds that $S(r,t,x(\textbf{A},\delta))\leq\textsf{park}(r,t)$. Thus, we conclude that $x(\textbf{A},\delta)\in X$.